### v1

@@ 4.7.5 Implementation Details
We adopted a loosely coupled architecture in which the Django-powered interface and the AEA-powered agent network were developed independently. These two components were integrated into a unified system through the use of RESTful APIs and the WebSocket protocol. In total, we implemented seven distinct agents within the agent network, comprising one supplier, one wholesaler, two retailers, one logistics, two 3PL providers, and one admin. These agents were individually developed using the AEA framework and collaborate by exchanging messages following predefined interaction protocols.
@@
The AEA framework provides a base structure for agents, resulting in the directory structure employed for agent development in this prototype. The agent directory structure is illustrated in Figure 11. As shown, an agent is composed of four main packages for further development: connections, contracts, protocols, and skills. These packages are responsible for implementing agent-specific business logic. In out implementation, we used built-in implementations for connections, and protocols. Since this prototype does not include any features related to blockchain, we did not consider contracts, which typically contains wrappers for smart contract logic. Our primary focus was on skills, which define the specific services and capabilities of an agent. An agent may possess multiple skills, each tailored to handle interactions with other agents under specific scenarios. Skills are divided into types of actions: reactive and proactive, implemented through a set of handlers and behaviours, respectively. Handlers are responsible for processing incoming messages following a specified protocol, constituting an agent’s reactive behaviour. Behaviours encompass self-initiated actions driven by internal logic rather than external messages. Additionally, a skill incorporates data models for maintaining shared state within a skill, including description of the services that an agent with this skill can offer.
@@
To manage CPU-bound or long-running executions, the framework provides the capability to create tasks within handlers and behaviours. We utilised this capability to handle long-executing tasks such as monitoring the delivery process. Decision-making logic, such as assessing a proposal, is organised into a strategy file within a skill. Additionally, a skill may contain files related to agent dialogue management (dialogues), database connections (db_communications), and miscellaneous utilities (helpers). An agent’s capabilities are configurable through YAML files, allowing tasks such as configuring a skill or specifying a RESTful service. For a more comprehensive understanding of the AEA framework, refer to the work by Minarsch et al. [2021].
@@
The implemented agents have access two types of data sources: databases and IoT sensors. Agents such as supplier, wholesaler, and retailer manage their own inventory systems, which were simulated by updating their respective inventory table when products are inbound or outbound. PostgreSQL was used as the database system to create these tables, which contains a list of products and their essential details, such as product name and quantity. The 3PL agents required access to IoT sensors to for delivery monitoring. This monitoring process was simulated by periodically read the data that was pre-captured using the method described in Section 4.7.4.
@@
The web interface for this prototype was built using Django and its plugin Channels, as the underlying technologies. WebSockets and RESTful APIs were used to exchange data and control between the web interface and the agent network, effectively connecting the frontend and backend of the system. Real-time data, including vehicle location, agent messages, and system notifications, were transmitted to the interface using WebSockets, in which consumers was created to handle these data. These data were then rendered using appropriate frontend techniques. For instance, Chart.js was used for visualising IoT data, while OSM and leaflet were employed for tracking and monitoring the delivery vehicle. The agent chat room was developed from scratch without relying on specific libraries. The ordering panel was implemented using Ajax, a set of client-side techniques that enable web applications to send and receive from a server (i.e., an agent) asynchronously without interfering with the display and behaviour of the existing page. To enhance the layout and styling of the interface, Bootstrap and JQuery were utilised. Further details about the tools used in developing the prototype are provided in Section 4.7.1.
4.8 Showcase
Figure 12: A screenshot of the startup interface of the prototype. This interface consists of four areas: the top left panel (ordering), the top middle panel (map for delivery monitoring), the right panel (agent chat room), and the bottom panel (IoT data visualisation). All of areas are is in their initial states, with empty content or default values.
Once the OEF, all agents, the Django built-in web server, and the PostgreSQL service are started, the prototype is ready for use. Upon initialisation, each agent will execute a registration behaviour. This behaviour is a one-shot proactive action that involves sending a message to the OEF for registering their service. This registration allows them to be discovered by other agents in the system.
The prototype can be accessed via a web browser. Figure 12 shows the startup interface of the prototype, which includes the four panels as described in Section 4.7.3. Initially, no data is displayed on the interface besides the default values in the ordering panel. These default values include the default quantity and unit price of the meat to purchase, default performance and location requirements for vendors, and the default scenario. In this section, we will showcase this A2SC prototype, focusing on the two processes: procurement and replenishment.
At the beginning of the execution,the wholesaler’s inventory is empty. To prevent running out of stock when the retailer procures meat from the wholesaler, we initiate a replenishment process. We select the ‘replenish’ scenario and click the ‘Launch’ button with other default values. Clicking the ‘Launch’ button triggers the replenishment behaviour of the wholesaler agent. The wholesaler agent first establish connection with the OEF and sends it a query to search suitable suppliers. In response, the OEF returns a message containing a list of supplier agents that meat the query conditions back to the sender. This process is essentially a matchmaking step, matching a service requester with a suitable provider. The wholesaler can then start dialogues with the discovered suppliers, negotiating with a selected supplier for the purchase of products. The logistic agent also participates in this dialogue, and the supplier requests available delivery options from this agent. When all parties agree on the deal, the logistics agent randomly selects one 3PL from a set of eligible 3PL agents to fulfil the delivery. The delivery process begins, and the 3PL agent is responsible for monitoring and managing the delivery.
The resulting interfaces during the replenishment process are shown in Figures 12(a) and 12(b). Figure 12(a) shows the interface at the moment when dialogue is completed, and the delivery is just starting. At this stage, no IoT data has been collected yet, so there is no data displayed on the charts below the map. The notifications at the top right corner of the interface, as presented in Figure 12(a), provide key progress updates, such as indicating that the dialogue is done and the delivery has commenced. Figure 12(b) presents a screenshot of the interface during the delivery process. The delivery is visually represented by a moving vehicle with a blue-line trace on the map, along with an overlaid tracking number and a blinking red circle. The ambient condition inside the delivery vehicle are also captured and visualised on the charts.
When the product has been successfully delivered, the relevant data collected during the delivery journey are gathered, and a summary report is generated using a data exploration tool called Pandas profiling  [^15] https://pandas-profiling.ydata.ai/. This summary report can be accessed by clicking the ‘Report’ dropdown button or the stacked icons at the top right corner of the map. The aim of this summary report is to demonstrate the potential use of the data for evaluating the delivery service, and therefore, a quick exploratory data analysis tool was adopted for report generation. The corresponding interface screenshot is shown in Figure 12(c). A retailer’s automated procurement is similar to the wholesaler’s replenishment, as described in Section 4.2. When the procurement is complete, the entire trace of the delivery vehicle is shown on the map, as depicted in Figure 12(d).
The dialogues within this process occur through messages, which are displayed in the agent chat room area. Excerpts of agent chat messages are presented in Figures 14 As shown in Figures 14, messages have four properties: sender, recipient, performative, and message body (content). The message body contains data associated with the message, such as procurement orders, receipts, and delivery notifications. These data are represented using JSON (or Javascript Object Notation). JSON is a human-readable data interchange format suitable for storing and transmitting data in web applications with servers. A screen recording of this prototype, providing more detailed demonstrations, is accessible via this link  [^16] https://www.youtube.com/watch?v=WFcLAif-cHo.
5 Discussion and Limitations
The prototype provides a concrete system to showcase the implementation of an ASC using a MAS approach. However, it is worth noting that this prototype may not strictly qualify as a complete ASC. Instead, it is an experimental proof of concept aligned with the MISSI model — a conceptual model described in Xu et al. [2022]. In this implementation, the main is placed on the two central layers of the model: interconnection and integration. The MAS approach was adopted as the overarching architecture to facilitate the interconnection and integration of decentralised business entities along the supply chain, creating a cohesive whole. The coherence of this whole is achieved through interaction enabled by message flows among the participating agents. To facilitate effective interactions, we need to address the following three key aspects:
  • Representation: This involves formally describing the current status or knowledge of an issue, request, or response. It is about how we represent data, such as bids or events, in a structured and machine-readable manner.
  • Decision-making: This aspect concerns the evaluation of incoming requests and the process of making appropriate responses. It is about defining rules and criteria for determining the best course of action in response to various issues.
  • Social: The focus is on establishing shared standards of acceptable behaviour, protocols, and language that enable smooth interactions with others. It involves defining the norms and conventions that govern how agents communicate and collaborate within the system.
In this prototype, we have indeed considered these three aspects, but our implementation is relatively basic and limited in scope. For representation, we primarily addressed how to represent the service that agents can offer and some data objects in the scenario, such as purchase orders and receipts. Regarding decision-making, our approach was rule-based and relatively simplistic. For example, the strategy of a wholesaler for accepting or rejecting a purchase order relies on a simple rule: it accepts the order if it has sufficient product to supply and the purchase price is higher than expected; otherwise, it rejects the order. We utilised the contract-net protocol to govern interactions and coordinate task allocation. However, it is important to note that this protocol alone may not suffice for handling complex real-world interactions that involve intricate negotiations. In terms of social interactions, our prototype does not incorporate formal languages for knowledge representation and social protocols. This prototype represents an initial step toward building a coherent A2SC system, which mainly focuses on automating a series of functions through data exchange. We underscore the need for carefully designed formal or even natural languages for knowledge representation and social protocols.
According to the autonomy manifold presented in Xu et al. [2022], this prototype falls within the automation-skewed region. However, both the intelligence and automation levels of the prototype are relatively low. The intelligence dimension primarily concerns decision-making capabilities. In this prototype, decision making is rudimentary, relying on a set of simple rules that consider only a few factors. To address real-world challenges in the supply chain domain, the agents need to be equipped with more sophisticated decision-making capabilities, enabling them to make informed decisions and propose appropriate courses of actions. The automation dimension of this prototype mainly focuses on automating data flow and process execution. This automation is facilitated through interaction among AEAs, which can request services or respond to requests, ultimately streamlining the overall process. However, this automation is limited to known data transmission paths, and the prototype neglects important aspects such as financial and product flows, which are integral to a functional supply chain. For an effective ASC system, it is crucial to address and enhance both the intelligence and automation dimensions to handle a broader range of scenarios and challenges.
The prototype includes two distinct processes: replenishment and wholesale. Both processes can be initiated manually by a user who clicks the ‘Launch’ button after completing an order form with the desired values. While the objective is to achieve level 2 autonomy, which implies process automation, the prototype falls short by only partially automating its functions. Automation is limited to ordering and order fulfilment in a controlled simulation environment, with financial and physical flows remaining unrealised. Additionally, SCM functions such as planning are absent from this prototype, which are essential for an ASC reach higher autonomy levels.
This prototype involves approximately ten stakeholders, which is significantly fewer than what is typically encountered in real-world supply chain scenarios. Managing a large number of agents can introduce complexities in various aspects, including architectural scalability, protocol design, and managerial challenges. In reality, ASCs must address these complexities to create a dynamic and adaptable society of living agents capable of self-adjusting to accommodate changes and uncertainties.
Furthermore, as highlighted in Xu et al. [2022], the development of ASCs involves the consideration of critical aspects that extend beyond technological efforts. These aspects include cyber security, data security and privacy, trustworthiness, and platform neutrality. Effectively addressing these dimensions need collaborative efforts from all participants within the ASC ecosystem. As an initial proof-of-concept, this implementation may not have fully addressed all of these aspects. However, its adoption of an MAS approach for integrating distributed entites represents a promising step toward addressing platform neutrality in the broader ASC domain.
6 Conclusion and Future Work
This paper focuses on the implementation of an autonomous supply chain (ASC) system, which is an emerging concept within the field of SCM. Specifically, we introduce a model for ASC implementation known as the agent-based autonomous supply chain (A2SC). A2SC is an ASC system that is realised through the utilisation of a multi-agent system (MAS) approach.
We have presented a methodology designed for the analysis and design of A2SC systems, which is adapted from a classic agent-oriented analysis and design framework, the Gaia methodology [Wooldridge et al., 2000]. We have provided a detailed concrete case study, the autonomous meat supply chain, demonstrating the practical implementation of A2SC systems using our proposed methodology. Additionally, we presented a system architecture and a toolkit for developing A2SCs. While the architecture and tooklit are specifically designed for this case, they can offer valuable guidelines for realising other similar decentralised systems.
As discussed in the previous section, this work represents an initial attempt to implement an ASC system, and it comes with certain limitations. Our future endeavours will primarily focus on addressing these limitations, which include 1) development of appropriate languages for agent communication and content representation within A2SCs, 2) designing or introducing protocols capable of handling more complex scenarios, such as bargaining and multi-issue negotiation, 3) scaling up the problem, including expanding the number of stakeholders, and 4) conducting experiments in less restricted setting. Furthermore, we plan to explore the potential applications of emerging technologies and concepts within ASCs, such as Digital Twin (DT) and artificial general intelligence (AGI). A2SCs can be viewed as a unique form of supply chain digital twins, serving as not only virtual representations but also control centres for actual supply chains. This perspective positions DT technologies and related concepts as promising avenues for advancing ASC development. Additionally, the recent tremendous progress witnessed in AGI systems, particularly in the domain of Large Language Models such as ChatGPT [OpenAI, 2023] and Llama [Touvron et al., 2023], offer intriguing opportunities for A2SCs. These advancements enable the creation of intelligent agents capable of making human-like decisions — a pivotal capability for effectively managing various supply chain tasks in A2SCs. We will thus explore the potential of AGI technologies in development of ASC systems.
In conclusion, this study, to the best of our knowledge, is the first work to report on a technical implementation case of an ASC system. It demonstrates a promising pathway for the development of an effective ASC system. The methodology, system architecture, toolkit for implementation, and agent roles and services model presented in this paper shed light on the future research and technical implementation of ASC systems in various scenarios.
Acknowledgement
This work was funded by the Research England’s Connecting Capability Fund (grant number: CCf18-7157): Promoting the Internet of Things via Collaboration between HEIs and Industry (Pitch-In) and the EPSRC Connected Everything Network Plus under grant EP/S036113/1.
Appendix A Additional Figures

### v2

@@ 4.7.5 Implementation Details
We adopted a loosely coupled architecture, in which the frontend interface and the backend agent network can be developed independently. These two components were integrated into a system through the use of RESTful APIs and the WebSocket protocol. In total, we implemented seven distinct agents within the agent network, comprising one supplier, one wholesaler, two retailers, one logistics, two 3PL providers, and one admin. These agents were individually developed using the AEA framework and can collaborate by exchanging messages by following predefined interaction protocols. More details about implementation can be found in B.
4.8 Showcase
Figure 11: A screenshot of the startup interface of the prototype. This interface consists of four areas: the top left panel (ordering), the top middle panel (map for delivery monitoring), the right panel (agent chat room), and the bottom panel (IoT data visualisation). All of areas are in their initial states, containing empty content or default values.
Once the OEF, all agents, the web server, and the database are started up, the prototype is ready for running. Upon initialisation, each agent performs a registration behaviour — a one-shot proactive action that involves sending a message to the OEF for registering their service. This registration enables their discovery by other agents in the system.
The prototype can be accessed via a web browser. Figure 11 shows the startup interface with four panels as described in Section  4.7.3. Initially, the interface shows no data except for default values in the ordering panel. These default values include the default quantity and unit price of the meat for purchase, default performance and location requirements for vendors, and the default scenario. This section showcases this A2SC prototype, focusing on the two main processes: procurement and replenishment.
At the start of execution, the wholesaler’s inventory is empty. To prevent out of stock when the retailer procures meat from the wholesaler, we first start a replenishment process by selecting the ‘replenish’ scenario and clicking the ‘Launch’ button with default values. Clicking ‘Launch’ button triggers the replenishment behaviour of the wholesaler agent. The wholesaler agent first establish connection with the OEF and sends a query to search suitable suppliers. The OEF responds with a message listing supplier agents that meat the query conditions. This process is essentially a matchmaking step, matching a service requester with a suitable provider. The wholesaler then engages in dialogues with discovered suppliers, negotiating with a selected supplier for purchasing products. The logistic agent participates in this dialogue, and the supplier requests available delivery options from this agent. Once all parties agree, the logistics agent randomly selects one of eligible 3PL agents to fulfil the delivery. The delivery process begins; and the 3PL agent is responsible for monitoring and managing this process.
The resulting interfaces during the replenishment process are shown in Figures 13(a) and 13(b). Figure 13(a) shows the interface when the dialogue is completed, and the delivery is commencing. At this stage, no IoT data has been collected, and the charts below the map display no data. Notifications at the top right corner, as presented in Figure 13(a), provide key progress updates, such as the completion of the dialogue and the commence of delivery. Figure 13(b) presents the interface during the delivery process, featuring a moving vehicle on the map with a blue-line trace, an overlaid tracking number, and a blinking red circle. Ambient condition inside the delivery vehicle is also captured and visualised on the charts.
When the product has been successfully delivered, data collected during the delivery journey are utilised to generate a summary report using a data profiling tool. Access to this report is available through the ‘Report’ dropdown button or the stacked icons at the top right corner of the map. The purpose of this report is to demonstrate the data’s potential for evaluating the delivery service. The corresponding interface is shown in Figure 13(c). The automated procurement for a retailer is similar to the wholesaler’s replenishment, as described in Section  4.2. When the procurement is complete, the entire trace of the delivery vehicle is displayed on the map, as shown in Figure 13(d).
The dialogues in this process take place through messages, which are displayed in the agent chat room area. Excerpts of agent chat messages are provided in Figure 14. These messages have four properties: sender, recipient, performative, and message body (content). The message body contains data (represented using JSON) associated with the message, such as procurement orders, receipts, and delivery notifications. A more detailed demonstration, including a screen recording of this prototype, is available via this link  [^7] https://www.youtube.com/watch?v=WFcLAif-cHo.
5 Discussion and Limitations
The prototype provides a concrete example to demonstrate the implementation of an ASC using an MAS approach. The prototype comprises two distinct integrated processes: replenishment and wholesale. Both processes can be manually initiated by a user clicking the ‘Launch’ button after filling out an order form, or they can be set up to start automatically. In addition to order placement, these two processes involve a set of automated functions, including supplier selection, negotiation, logistic service selection, and delivery monitoring and evaluation, inventory management. These functions are orchestrated into automate processes through messaging between agents and facilitate the management of the “virtual” product flow and its associated information flow. While this work demonstrates a promising approach to achieving ASC, as one of the earliest technical implementations, it also comes with some limitations.
The scenario adopted in the prototype is a simplified version of real-world operations in a meat supply chain. Its automation is mainly limited to the ordering and order fulfilment processes in a controlled simulation environment. While targeting level 2 autonomy, which requires process automation according to Xu et al. (2023b), the prototype only achieves partial process automation. It primarily addresses the management of the information flow while digitally representing the physical flow and disregarding the financial flow. This simplification reduces the complexity of the ASC system analysis and design, however, it may also widen the gap between this work and realistic systems.
While this prototype successfully integrates automated functions and connects decision making points, it lacks the ability to reconfigure and adapt to disruptions, which is essential for ASCs to achieve resilience. However, the proposed approach lays the foundation to facilitate this capability by enhancing agents with more advanced decision making skills. Realising this capability will be one of our primary future endeavours. Additionally, crucial SCM functions such as planning are absent, hindering the prototype from reaching higher autonomy levels.
This prototype involves around ten stakeholders, which is significantly fewer than what is typically encountered in real-world supply chain scenarios. Managing larger number of agents may introduce complexities in architectural scalability, protocol design, and managerial challenges etc. In real world applications, ASCs must address these complexities to create an open, dynamic, and adaptive society of autonomous agents that can self-adjust to accommodate changes and uncertainties.
Moreover, this prototype may not fully realise a complete ASC automated processes for the three flows. Instead, it represents an experimental proof of concept, implementing the MISSI model, a conceptual model described in Xu et al. (2023b). This prototype mainly focuses on the two central layers of the MISSI model: interconnection and integration. It adopts the MAS approach as the overarching architecture, facilitating the interconnection and integration of decentralised business entities across the supply chain, thereby creating a cohesive system. The coherence of this system relies on interactions through messaging between participating agents. To achieve effective interactions, three key aspects need attention:
  • Representation: This aspect involves formally describing the current status or knowledge of an issue, request, or response. It pertains to how data, such as bids or events, is structured and made machine-readable.
  • Decision-making: This aspect concerns the evaluation of incoming requests and the process of making appropriate responses, which involves defining rules and criteria for determining the best course of actions in response to various issues.
  • Social: This aspect focuses on establishing shared standards of acceptable behaviour, protocols, and languages that facilitate smooth interactions. It involves defining the norms and conventions that govern agent communication and collaboration within the system.
In this prototype, we have addressed these three aspects — representation, decision-making, and social interactions. However, our implementation is relatively basic and has limitations. For representation, we focused on how agents represent their services and certain data objects, such as purchase orders and receipts. Decision-making is rule-based and relatively straightforward. For instance, a wholesaler’s strategy for order acceptance is based on a simple rule: accepts the order if it has sufficient stock to supply and the purchase price is higher than expected; otherwise, it rejects the order. We employed the contract-net protocol for regulating interactions and task coordination, However, this protocol alone may not be adequate for handling complex real-world interactions. Socially, the prototype lacks formal languages for knowledge representation and social protocols. We underscore the need for well-designed formal or natural languages for knowledge representation and social protocols.
According to the autonomy manifold presented in Xu et al. (2023b)), this prototype falls within the automation-skewed region. However, both the intelligence and automation levels of the prototype are still relatively low. Its intelligence dimension, primarily focusing on decision-making capabilities, is basic, relying on simple rules that consider only a few main factors. To address real-world supply chain challenges, agents need more sophisticated decision-making capabilities, enabling them to make informed decisions and propose appropriate courses of actions. Its automation dimension mainly focuses on automating data flow and process execution through agent interactions, albeit limited to known data paths. However, the prototype overlooks crucial aspects such as financial and product flows, which are integral to a functional supply chain. For an effective ASC system, both intelligence and automation dimensions must be addressed, thereby enabling to handle a broader range of scenarios and challenges.
Meanwhile, as emphasised in Xu et al. (2023b), the development of ASCs requires consideration of critical aspects extending beyond technological efforts. These aspects include cyber security, data security and privacy, trustworthiness, and platform neutrality. Effectively addressing these aspects need collaborative efforts from all participants within the ASC ecosystem. While this initial proof-of-concept prototype may not fully addressed all of these aspects, its adoption of an MAS approach for integrating distributed entities represents a promising step toward addressing platform neutrality in the broader ASC domain.
6 Conclusion and Future Work
This paper focuses on realising an autonomous supply chain (ASC) system, an recently emerging concept in the domain of SCM. Specifically, it introduces the agent-based autonomous supply chain (A2SC) through a multi-agent system (MAS) approach.
We have presented a methodology for A2SC analysis and design, which is adapted from a classic agent-oriented analysis and design framework, the Gaia methodology (Wooldridge et al., 2000). We have provided a detailed concrete case study, the autonomous meat supply chain, demonstrating the technical implementation of an A2SC system using the MAS approach. Additionally, we presented a system architecture and a toolkit for developing A2SCs, offering guidelines for implementing similar decentralised systems.
This work marks an initial attempt to implement an ASC system, but there are certain limitations, as discussed in the previous section. Future efforts will mainly focusing on addressing these limitations, which include 1) development of appropriate languages for agent communication and content representation for A2SC; 2) designing or introducing protocols capable of handling more complex scenarios, such as bargaining and multi-issue negotiations; 3) scaling up the problem, including expanding the number of stakeholders; 4) conducting experiments in less restricted setting; and 5) equipping agents with intelligent decision making and planning capabilities. Furthermore, we will explore the potential applications of emerging technologies and concepts within ASCs, such as Digital Twin (DT) (Sharma et al., 2022) and large language models (LLMs) (Bommasani et al., 2021; Xu et al., 2023a). A2SCs can be viewed as a unique form of supply chain digital twins, which are not only virtual representations but also control centres for actual supply chains. This perspective positions DT technologies and related concepts as promising avenues for advancing ASC development. Recent progress in LLMs, such as ChatGPT (OpenAI, 2023) and Llama (Touvron et al., 2023), may enable the creation of AI agents that are truly capable of making decisions and planning on a more broad range of supply chain tasks with minimal human intervention.
In conclusion, this study is among the earliest explorations of implementing an ASC system with automated integrated processes by following a software engineering design methodology. It demonstrates a promising pathway for effective ASC system development. The methodology for A2SC analysis and design, system architecture, toolkit for implementation, and agent roles and services model presented in this paper shed light on the future research and technical implementation of ASC systems in various scenarios.
Acknowledgement
This work was supported by the Research England’s Connecting Capability Fund (grant number: CCF18-7157): Promoting the Internet of Things via Collaboration between HEIs and Industry (Pitch-In) and the EPSRC Connected Everything Network Plus under grant EP/S036113/1.
Appendix A Toolkit for Implementation
The tools and frameworks that were employed for developing the A2SC prototype are described below:
  • AEA Framework: The AEA (Autonomous Economic Agent) framework (Minarsch et al., 2021) from Fetch.ai  [^8] https://fetch.ai/ technology stack is a Python-based toolkit for constructing autonomous agents known as AEAs. These AEAs operate independently on behalf of their owners and autonomously execute actions to achieve predefined objectives. We used this framework to develop the agent network in the prototype.
  • OEF (Open Economic Framework): The OEF serves as a decentralised communication, search and discovery system for AEAs, providing protocols, languages, connection mechanisms that facilitate agents search, interaction discovery, and engagement within established social norms. It establishes a virtual environment where the inhabitants, AEAs, can operate and interact. In Fetch.ai technology stack, the AEA framework and the OEF represent the top two layers, positioned above the smart ledger layer. For this implementation, we exclusively utilised the AEA framework and the OEF to develop the prototype, as they are capable of functioning independently without the smart ledger.
  • REST and OpenAPI: REST (REpresentational State Transfer) Fielding (2000) is an architectural style for providing interface standards between decoupled components on the web. RESTful APIs are typically based on HTTP methods to access resources via URL-encoded parameters and the use of JSON or XML for data transmission. All agents in the prototype were equipped with RESTful APIs, enabling them communicate with each other without the OEF and access external resources. However, these agents need to expose their APIs to other in a standard manner. The OpenAPI specification (OAS) (OpenAPI Initiative, 2021) was therefore adopted for describing these interfaces. The OAS defines a standard, programming language-agnostic interface description for RESTful APIs, allowing both humans and machines to discover and comprehend a web service without requiring access to source code and extra documentation. In this prototype, we define interface description files (.yaml) by adhering to the OAS.
  • WebSockets: WebSockets, which enable bi-directional, full-duplex communication, are well-suited for frequently updated or real-time applications. We thus used WebSockets (Fette and Melnikov, 2011) for services such as notification, monitoring, and agent chatting. The above tools are relevant to backend development and architectural design. We will describe the tools/frameworks used for frontend development in this following paragraphs.
In addition to these major tools, we incorporated complementary tools and frameworks such as Bootstrap  [^14] https://getbootstrap.com/, JQuery  [^15] https://jquery.com/, and Ajax.
Appendix B Implementation Details
We adopted a loosely coupled architecture, in which the frontend interface and the backend agent network can be developed independently. These two components were integrated into a unified system through the use of RESTful APIs and the WebSocket protocol. In total, we implemented seven distinct agents within the agent network, comprising one supplier, one wholesaler, two retailers, one logistics, two 3PL providers, and one admin. These agents were individually developed using the AEA framework and collaborate by exchanging messages following predefined interaction protocols.
@@ Appendix B Implementation Details
The AEA framework provides a base structure for agents, resulting in the directory structure employed for agent development in this prototype. The agent directory structure is illustrated in Figure 12. An agent directory is composed of four main packages for further development: connections, contracts, protocols, and skills. These packages are responsible for implementing agent-specific business logic. In out implementation, we used built-in implementations for connections, and protocols. Since this prototype does not include any features related to blockchain, we did not consider contracts, which typically contains wrappers for smart contract logic. Our primary focus was on skills, which define the specific services and capabilities of an agent. An agent may possess multiple skills, each tailored to handle interactions with other agents under specific scenarios. Skills are divided into types of actions: reactive and proactive, implemented through a set of handlers and behaviours, respectively. Handlers are responsible for processing incoming messages following a specified protocol, constituting an agent’s reactive behaviour. Behaviours encompass self-initiated actions driven by internal logic rather than external messages. Additionally, a skill incorporates data models for maintaining shared state within a skill, including description of the services that an agent with this skill can offer.
@@
To manage CPU-bound or long-running executions, the framework provides the capability to create tasks within handlers and behaviours. We utilised this capability to handle long-executing tasks such as monitoring the delivery process. Decision-making logic, such as assessing a proposal, is organised into a strategy file within a skill. A skill may contain files related to agent dialogue management (dialogues), database connections (db_communications), and miscellaneous utilities (helpers). An agent’s capabilities are configurable, allowing tasks such as configuring a skill or specifying a RESTful service. Refer to the work by Minarsch et al. (2021) for more details on the AEA framework.
@@
The implemented agents can access two types of data sources: databases and sensors. Agents such as supplier, wholesaler, and retailer manage their own inventory systems, which were simulated by updating their respective inventory table when products are inbound or outbound. The 3PL agents require access to sensors to for delivery monitoring, which was simulated by periodically read the data that was pre-captured using the method described in Section  4.7.4.
@@
The interface of this prototype was developed using Django and its plugin Channels. WebSockets and RESTful APIs were used to exchange data and control between the web interface and the agent network, effectively connecting the frontend and backend of the system. Real-time data, including vehicle location, agent messages, and system notifications, were transmitted to the interface using WebSockets, in which consumers was created to handle these data. These data were then rendered using appropriate frontend techniques. For instance, Chart.js was used for visualising IoT data, while OSM and leaflet were employed for tracking and monitoring the delivery vehicle. The agent chat room was developed from scratch without relying on specific libraries. The ordering panel was implemented using Ajax, a set of client-side techniques that enable web applications to send and receive from a server (i.e., an agent) asynchronously without interfering with the display and behaviour of the existing page. To enhance the layout and styling of the interface, Bootstrap and JQuery were utilised. Further details about the tools used in developing the prototype are provided in Section  4.7.1.
Appendix C Additional Figures